OpenStudy (anonymous):

Which equation correctly states Coulomb's law? I tried researching but the answers on my quiz don't match the equations I've searched.

OpenStudy (anonymous):

My choices are: Force = k(q1q2)/d2 Force = k(q2)/d Force = k(d1d2)/q